1. The savings rate in a country is 25%. The government wants to expand GDP by $300 Billion. How much must they increase Government spending to reach that goal?
Answer #1

Answer 1

It is given that saving rate is 25%. Thus Marginal Propensity to save = 0.25.

Multiplier = 1/MPS = 1/0.25 = 4.

Thus $1 increase in autonomous expenditure(like government spending) will increase GDP by $4. Here we want to increase GDP by 300 billion. Thus in order to increase GDP by 300 billion Government has to increase government spending by 300 billion/4 = 75 billion.

Hence, they must increase Government spending by $75 billion to reach that goal.

Note In the above problem we have assumed that MPI = Marginal Propensity to import = 0 that tax rate = 0 and if there are taxes then those are Lump sum tax.